# Vanishing criteria for Ceresa cycles

### v1

@@ 1.4. Structure of paper
In §2 we collect some standard results on Chow groups, Chow motives and Ceresa cycles. In §3 we prove the cohomological vanishing criteria (Theorems A and B) and give some examples. In §4 we study the family of Picard curves in detail and prove Theorem C. Finally, in §5 we introduce the Ceresa vanishing loci $V_{g}^{\mathrm{rat}},V_{g}^{\mathrm{alg}}\subset\mathcal{M}_{g}$ and determine which automorphism strata they contain in genus $3$, proving Theorem D.
@@ 1.5. Acknowledgements
We thank Jeffrey Achter, Ben Moonen, Burt Totaro and Congling Qiu for helpful conversations. This research was carried out while the first author was a Research Fellow at St John’s College, University of Cambridge, which he thanks for providing excellent working conditions. The second author was funded by the European Research Council (ERC, CurveArithmetic, 101078157).
@@ 2.1. Chow groups
Let $k$ be field. A variety is by definition a separated scheme of finite type over $k$. We say a variety is nice if it is smooth, projective and geometrically integral. If $X$ is a smooth and geometrically integral variety and $p\in\{0,\dots,\dim X\}$, let ${\rm CH}^{p}(X)$ denote the Chow group (with ${\rm\mathbb{Z}}$-coefficients) of codimension $p$ cycles modulo rational equivalence. If $Z\subset X$ is a closed subscheme of codimension $p$, we denote its class in ${\rm CH}^{p}(X)$ by $[Z]$ (using [44, Tag 02QS] if $Z$ is not integral).
@@ 2.2. Base change, specialization and families
To discuss specialization in the next lemma, let $R$ be a discrete valuation ring with fraction field $K$ and residue field $k$. Let $X\rightarrow{\rm Spec}(R)$ be a smooth, projective morphism with geometrically integral fibers, so the generic and special fibers $X_{K}$ and $X_{k}$ are nice varieties over $K$ and $k$ respectively. In this setting, Fulton has defined [20, §20.3] a specialization morphism $\mathrm{sp}\colon{\rm CH}^{p}(X_{K})\rightarrow{\rm CH}^{p}(X_{k})$ for every $0\leq p\leq\dim(X_{K})$. It has the property that if $Z\subset X$ is a closed integral subscheme of codimension $p$, flat over $R$, then $\mathrm{sp}([Z_{K}])=[Z_{k}]$.
@@
Let $\mathcal{C}\rightarrow{\rm Spec}(R)$ be an integral, projective, flat, proper and regular model of $C$, which exists by [30, Proposition 10.1.8]. Let $\mathrm{Hilb}_{X/R}\rightarrow{\rm Spec}(R)$ be the Hilbert scheme of the projective morphism $X\rightarrow{\rm Spec}(R)$. The closed subscheme $Z$ determines a $K$-morphism $\alpha\colon C\rightarrow\mathrm{Hilb}_{X/R}$. Since $C$ is connected, the image of $\alpha$ is contained in an open and closed subscheme $\mathrm{Hilb}_{X/R}^{\phi}$ with fixed Hilbert polynomial. The surface $\mathcal{C}$ is regular, $\mathrm{Hilb}^{\phi}_{X/R}\rightarrow{\rm Spec}(R)$ is projective [3, Corollary (2.8)] and we may view $\alpha$ as a rational map $\mathcal{C}\dashrightarrow\mathrm{Hilb}_{X/R}^{\phi}$. Therefore by [30, Theorem 9.2.7] there exists a birational map $\widetilde{\mathcal{C}}\rightarrow\mathcal{C}$, obtained by successively blowing up closed points in the special fiber, and an $R$-morphism $\tilde{\alpha}\colon\widetilde{\mathcal{C}}\rightarrow\mathrm{Hilb}^{\phi}_{X/R}$ with generic fiber $\alpha$. In other words, there exists a closed subscheme $\mathcal{Z}\subset\widetilde{\mathcal{C}}\times_{R}X$, flat over $\widetilde{\mathcal{C}}$, whose generic fiber equals $Z$.
@@ 2.3. Chow motives
We recall a few relevant facts about the category $\mathsf{Mot}(k)$ of (pure, contravariant) Chow motives $M$ over $k$ with ${\rm\mathbb{Q}}$-coefficients; see [42] or [38, §2] for basic definitions.
Denote the Lefschetz motive by $\mathbb{L}$, its $n$th tensor power by $\mathbb{L}^{n}$ and write $M(n)=M\otimes\mathbb{L}^{n}$. Following [38, §2.5], the Chow group in codimension $p$ of $M$ is by definition ${\rm CH}^{p}(M)={\rm Hom}_{\mathsf{Mot}(k)}(\mathbb{L}^{p},M)$. If $M=\mathfrak{h}(X)$ where $X$ is smooth projective over $k$, then ${\rm CH}^{p}(M)={\rm CH}^{p}(X)_{{\rm\mathbb{Q}}}$. (Beware that we need to take ${\rm\mathbb{Q}}$-coefficients on the right hand side.) If $C$ is a nice curve over $k$ and $\varphi\colon\mathfrak{h}(C)\rightarrow M(p-1)$ a morphism in $\mathsf{Mot}(k)$, we obtain a homomorphism of abelian groups ${\rm CH}^{1}(\varphi)\colon{\rm CH}^{1}(C)_{{\rm\mathbb{Q}}}\rightarrow{\rm CH}^{p}(M)$. We define ${\rm CH}^{p}(M)_{\mathrm{alg}}$ to be the union of the images of ${\rm CH}^{1}(\varphi)$ ranging over all pairs $(C,\varphi)$ as above. This coincides with ${\rm CH}^{p}(X)_{\mathrm{alg}}\otimes{\rm\mathbb{Q}}$ when $M=\mathfrak{h}(X)$, see [2, Corollary 3.13].
@@ 2.4. Chow–Künneth decomposition for abelian varieties
Let $A/k$ be a $g$-dimensional abelian variety. Deninger and Murre [18, §3] have constructed a canonical Chow–Künneth decomposition
@@ 2.5. The Lefschetz decomposition for abelian varieties
Let $A/k$ be an abelian variety with polarization $\lambda\colon A\rightarrow A^{\vee}$. Let $\ell\in{\rm CH}^{1}_{(0)}(A)$ be the unique class satisfying $2\ell=(1,\lambda)^{*}\mathcal{P}$, where $\mathcal{P}\in{\rm Pic}(A\times A^{\vee})$ is the Poincaré bundle. Künnemann has shown [26, Theorem 5.2] that intersecting with $\ell^{g-i}$ induces an isomorphism $\mathfrak{h}^{i}(A)\rightarrow\mathfrak{h}^{2g-i}(A)(g-i)$ for all $0\leq i\leq g$. This induces a Lefschetz decomposition of the Chow–Künneth components $\mathfrak{h}^{i}(A)$, see [26, Theorem 5.1]. We are chiefly interested in the components $\mathfrak{h}^{3}(A)$ and $\mathfrak{h}^{2g-3}(A)$ when $g\geq 2$; in that case the Lefschetz decomposition has the form $\mathfrak{h}^{3}(A)=\mathfrak{h}^{3}_{\mathrm{prim}}(A)\oplus\ell\cdot\mathfrak{h}^{1}(A)$ and $\mathfrak{h}^{2g-3}(A)=\mathfrak{h}^{2g-3}_{\mathrm{prim}}(A)\oplus\ell^{g-2}\cdot\mathfrak{h}^{1}(A)$. It has the property that the isomorphism $\ell^{g-3}\colon\mathfrak{h}^{3}(A)\rightarrow\mathfrak{h}^{2g-3}(A)(g-3)$ is a direct sum of isomorphisms $\mathfrak{h}^{3}_{\mathrm{prim}}(A)\rightarrow\mathfrak{h}^{2g-3}_{\mathrm{prim}}(A)$ and $\ell\cdot\mathfrak{h}^{1}(A)\rightarrow\ell^{g-2}\mathfrak{h}^{1}(A)$. Taking Chow groups, we get a decomposition ${\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(A))={\rm CH}^{g-1}(\mathfrak{h}_{\mathrm{prim}}^{2g-3}(A))\oplus{\rm CH}^{g-1}(\ell^{g-2}\cdot\mathfrak{h}^{1}(A))$.
@@ 2.6. Beauville components of $C$
(Here we view $[0]-[x_{e}]$ as an element of ${\rm CH}_{0}(J)$.)
@@
By definition and the discussion in §2.4, $[C]_{1}\in{\rm CH}_{(1)}^{g-1}(J)={\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J))$. Since $x_{e}$ is torsion, $(n)_{*}([0]-[x_{e}])=0$ for some integer $n\geq 1$. The decomposition (2.2) implies that $[0]=[x_{e}]$ in ${\rm CH}_{0}(J)_{{\rm\mathbb{Q}}}$. Therefore (2.3) shows that $[C]_{0}^{\star(g-2)}\star[C]_{1}=0$. Using properties of the $\mathfrak{sl}_{2}$-action on ${\rm CH}^{*}(J)_{{\rm\mathbb{Q}}}$ (in the sense of [33, §1.3]), this implies that $\ell\cdot[C]_{1}=0$. Since ${\rm CH}^{g-1}(\mathfrak{h}^{2g-3}_{\mathrm{prim}}(J))$ equals the kernel of $\ell\cdot(-)\colon{\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J))\rightarrow{\rm CH}^{g-1}(\mathfrak{h}^{2g-1}(J)(1))$, the corollary follows. ∎
@@ 2.7. Ceresa cycles
Using the expression (2.4), it suffices to prove $[C]_{1}=0$ implies $[C]_{s}=0$ for all $s\geq 1$. This follows from the third centered equation of [46, Theorem 1.5.5]. The proof for $\bar{\kappa}(C)$ is identical. ∎
@@ 3. Vanishing criteria in the Chow and Griffiths groups
By the hypothesis and the Lefschetz isomorphism, $\operatorname{H}^{*}(\mathfrak{h}_{\mathrm{prim}}^{2g-3}(J)^{G})=\operatorname{H}^{2g-3}_{\mathrm{prim}}(J)^{G}=0$. Since any summand of the motive of an abelian variety is finite-dimensional in the sense of Kimura [25, Example 9.1], it follows from Kimura’s [25, Corollary 7.3] that $\mathfrak{h}^{2g-3}_{\mathrm{prim}}(J)^{G}=0$ and hence $[C]_{1}=0$ in ${\rm CH}^{g-1}(J)_{{\rm\mathbb{Q}}}$. By Proposition 2.12, we conclude that $\kappa(C)=0$. ∎
@@
The genus $4$ curve $y^{3}=x^{5}+1$ satisfies $\operatorname{H}^{3}(J)^{{\rm Aut}(C)}=0$ [29, proof of Theorem 3.3], so $\kappa(C)=0$.
@@
We now show that the Hodge conjecture for $J\times A$ implies the claims of the final sentence. Fix mutually inverse isomorphisms of Hodge structures $\phi\colon\operatorname{H}^{3}(J)^{G}\rightarrow\operatorname{H}^{1}(A)(-1)$ and $\psi\colon\operatorname{H}^{1}(A)(-1)\rightarrow\operatorname{H}^{3}(J)$. By the Hodge conjecture, there exist morphisms of motives $\Phi\colon\mathfrak{h}^{3}(J)^{G}\rightarrow\mathfrak{h}^{1}(A)(-1)$ and $\Psi\colon\mathfrak{h}^{1}(A)(-1)\rightarrow\mathfrak{h}^{3}(J)^{G}$ (in other words, cycles on $J\times A$ with certain properties) with $\operatorname{H}^{*}(\Phi)=\phi$ and $\operatorname{H}^{*}(\Psi)=\psi$. Since $\mathfrak{h}^{1}(A)(-1)$ and $\mathfrak{h}^{3}(J)^{G}$ are Kimura finite-dimensional [25, Example 9.1] and $\operatorname{H}^{*}(\Phi)\circ\operatorname{H}^{*}(\Psi)$ and $\operatorname{H}^{*}(\Psi)\circ\operatorname{H}^{*}(\Phi)$ are the identity, it follows from [25, Proposition 7.2(ii)] (see also [4, Corollaire 3.16]) that $\Psi\circ\Phi$ and $\Phi\circ\Psi$ are themselves isomorphisms. Hence $\Phi$ and $\Psi$ are isomorphisms too and we conclude that $\mathfrak{h}^{3}(J)^{G}\simeq\mathfrak{h}^{1}(A)(-1)$.
Using the Lefschetz isomorphism $\mathfrak{h}^{2g-3}(J)\simeq\mathfrak{h}^{3}(J)(-g+3)$ of [26, Theorem 5.2], we obtain an isomorphism $\mathfrak{h}^{2g-3}(J)^{G}\simeq\mathfrak{h}^{1}(A)(-g+2)$. Similarly to the proof of Theorem A, we decompose $[\iota_{e}(C)]=\sum_{s=0}^{g-1}[C]_{s}$ with $[C]_{s}\in{\rm CH}^{g-1}_{(s)}(J)$ and observe that the class $[C]_{1}$ lies in ${\rm CH}_{(1)}^{g-1}(J)^{G}={\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J)^{G})$. Lemma 2.6 combined with the isomorphism $\mathfrak{h}^{2g-3}(J)^{G}\simeq\mathfrak{h}^{1}(A)(-g+2)$ shows that ${\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J)^{G})={\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J)^{G})_{\mathrm{alg}}$, hence every element of ${\rm CH}^{g-1}_{(1)}(J)^{G}$ lies in ${\rm CH}^{g-1}(J)_{\mathrm{alg},{\rm\mathbb{Q}}}$. Therefore $[C]_{1}\in{\rm CH}^{g-1}(J)_{\mathrm{alg},{\rm\mathbb{Q}}}$, hence the image of $[C]_{1}$ in ${\rm Gr}_{1}(J)_{{\rm\mathbb{Q}}}$ vanishes. Proposition 2.12 then implies that $\bar{\kappa}(C)$ vanishes too. ∎
@@
Let $C\colon y^{3}=x^{4}+ax^{2}+bx+c$ be a Picard curve. Then $\operatorname{H}^{0}(C,\Omega_{C}^{1})\simeq\chi\oplus\chi\oplus\chi^{2}$ as $C_{3}$-representations, where $\chi$ is a character of order $3$. It follows that $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{{\rm Aut}(C)}=0$, so that the condition of Theorem B is satisfied. We consider these curves in detail in the next section.
We exhibit two families of examples of Theorem B in genus $4$. It would be interesting to study these families in more detail (along the lines of what we do in the next section for Picard curves).
Example 3.5.
Consider for every $t\in{\rm\mathbb{C}}\setminus\{0,1,-1\}$ the $\mu_{5}$-cover of ${\rm\mathbb{P}}^{1}$ with equation
$C\colon y^{5}=x^{3}(x-1)^{2}(x-t).$
By [34, Lemma 2.7], the $\mu_{5}$-representation $\operatorname{H}^{0}(C,\Omega_{C})$ is the direct sum of the four nontrivial characters. It follows that $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{\mu_{5}}=0$ and $\operatorname{H}^{3}(J)^{\mu_{5}}$ is isomorphic to $\operatorname{H}^{1}(A)(-1)$, for some abelian surface $A$ (using Proposition 3.3). We have $D_{5}\subset{\rm Aut}(C)$ by [22, Corollary 1]. By the Kani-Rosen formula [23, Theorem B], the Jacobian $J$ is isogenous to a product of abelian surfaces of the form $\operatorname{Jac}(C/\tau)$, for some involution $\tau$. The Hodge conjecture in any codimension is known for products of abelian surfaces [40, Theorem 3.15], so $\bar{\kappa}(C)=0$ by Proposition 3.3.

### v2

@@ 1.4. Structure of paper
In §2 we collect some standard results on Chow groups, Chow motives and Ceresa cycles. In §3 we prove the cohomological vanishing criteria (Theorems A and B) and give some examples. In §4 we study the family of Picard curves in detail and prove Theorems C and D. Finally, in §5 we introduce the Ceresa vanishing loci $V_{g}^{\mathrm{rat}},V_{g}^{\mathrm{alg}}\subset\mathcal{M}_{g}$, prove Theorem D and determine which automorphism strata they contain in genus $3$, proving Theorem E.
@@ 1.5. Acknowledgements
We thank Jeffrey Achter, Irene Bouw, Edgar Costa, Ben Moonen, Burt Totaro and Congling Qiu for helpful conversations. This research was carried out while the first author was a Research Fellow at St John’s College, University of Cambridge, which he thanks for providing excellent working conditions. The second author was funded by the European Research Council (ERC, CurveArithmetic, 101078157).
@@ 2.1. Chow groups
Let $k$ be field. A variety is by definition a separated scheme of finite type over $k$. We say a variety is nice if it is smooth, projective and geometrically integral. If $X$ is a smooth and geometrically integral variety and $p\in\{0,\dots,\dim X\}$, let ${\rm CH}^{p}(X)$ denote the Chow group (with ${\rm\mathbb{Z}}$-coefficients) of codimension $p$ cycles modulo rational equivalence. If $Z\subset X$ is a closed subscheme of codimension $p$, we denote its class in ${\rm CH}^{p}(X)$ by $[Z]$ (using [49, Tag 02QS] if $Z$ is not integral).
@@ 2.2. Base change, specialization and families
To discuss specialization in the next lemma, let $R$ be a discrete valuation ring with fraction field $K$ and residue field $k$. Let $X\rightarrow{\rm Spec}(R)$ be a smooth, projective morphism with geometrically integral fibers, so the generic and special fibers $X_{K}$ and $X_{k}$ are nice varieties over $K$ and $k$ respectively. In this setting, Fulton has defined [23, §20.3] a specialization morphism $\mathrm{sp}\colon{\rm CH}^{p}(X_{K})\rightarrow{\rm CH}^{p}(X_{k})$ for every $0\leq p\leq\dim(X_{K})$. It has the property that if $Z\subset X$ is a closed integral subscheme of codimension $p$, flat over $R$, then $\mathrm{sp}([Z_{K}])=[Z_{k}]$.
@@
Let $\mathcal{C}\rightarrow{\rm Spec}(R)$ be an integral, projective, flat, proper and regular model of $C$, which exists by [34, Proposition 10.1.8]. Let $\mathrm{Hilb}_{X/R}\rightarrow{\rm Spec}(R)$ be the Hilbert scheme of the projective morphism $X\rightarrow{\rm Spec}(R)$. The closed subscheme $Z$ determines a $K$-morphism $\alpha\colon C\rightarrow\mathrm{Hilb}_{X/R}$. Since $C$ is connected, the image of $\alpha$ is contained in an open and closed subscheme $\mathrm{Hilb}_{X/R}^{\phi}$ with fixed Hilbert polynomial. The surface $\mathcal{C}$ is regular, $\mathrm{Hilb}^{\phi}_{X/R}\rightarrow{\rm Spec}(R)$ is projective [3, Corollary (2.8)] and we may view $\alpha$ as a rational map $\mathcal{C}\dashrightarrow\mathrm{Hilb}_{X/R}^{\phi}$. Therefore by [34, Theorem 9.2.7] there exists a birational map $\widetilde{\mathcal{C}}\rightarrow\mathcal{C}$, obtained by successively blowing up closed points in the special fiber, and an $R$-morphism $\tilde{\alpha}\colon\widetilde{\mathcal{C}}\rightarrow\mathrm{Hilb}^{\phi}_{X/R}$ with generic fiber $\alpha$. In other words, there exists a closed subscheme $\mathcal{Z}\subset\widetilde{\mathcal{C}}\times_{R}X$, flat over $\widetilde{\mathcal{C}}$, whose generic fiber equals $Z$.
@@ 2.3. Chow motives
We recall a few relevant facts about the category $\mathsf{Mot}(k)$ of (pure, contravariant) Chow motives $M$ over $k$ with ${\rm\mathbb{Q}}$-coefficients; see [47] or [42, §2] for basic definitions.
Denote the Lefschetz motive by $\mathbb{L}$, its $n$th tensor power by $\mathbb{L}^{n}$ and write $M(n)=M\otimes\mathbb{L}^{n}$. Following [42, §2.5], the Chow group in codimension $p$ of $M$ is by definition ${\rm CH}^{p}(M)={\rm Hom}_{\mathsf{Mot}(k)}(\mathbb{L}^{p},M)$. If $M=\mathfrak{h}(X)$ where $X$ is smooth projective over $k$, then ${\rm CH}^{p}(M)={\rm CH}^{p}(X)_{{\rm\mathbb{Q}}}$. (Beware that we need to take ${\rm\mathbb{Q}}$-coefficients on the right hand side.) If $C$ is a nice curve over $k$ and $\varphi\colon\mathfrak{h}(C)\rightarrow M(p-1)$ a morphism in $\mathsf{Mot}(k)$, we obtain a homomorphism of abelian groups ${\rm CH}^{1}(\varphi)\colon{\rm CH}^{1}(C)_{{\rm\mathbb{Q}}}\rightarrow{\rm CH}^{p}(M)$. We define ${\rm CH}^{p}(M)_{\mathrm{alg}}$ to be the union of the images of ${\rm CH}^{1}(\varphi)$ ranging over all pairs $(C,\varphi)$ as above. This coincides with ${\rm CH}^{p}(X)_{\mathrm{alg}}\otimes{\rm\mathbb{Q}}$ when $M=\mathfrak{h}(X)$, see [2, Corollary 3.13].
@@ 2.4. Chow–Künneth decomposition for abelian varieties
Let $A/k$ be a $g$-dimensional abelian variety. Deninger and Murre [19, §3] have constructed a canonical Chow–Künneth decomposition
@@ 2.5. The Lefschetz decomposition for abelian varieties
Let $A/k$ be an abelian variety with polarization $\lambda\colon A\rightarrow A^{\vee}$. Let $\ell\in{\rm CH}^{1}_{(0)}(A)$ be the unique class satisfying $2\ell=(1,\lambda)^{*}\mathcal{P}$, where $\mathcal{P}\in{\rm Pic}(A\times A^{\vee})$ is the Poincaré bundle. Künnemann has shown [30, Theorem 5.2] that intersecting with $\ell^{g-i}$ induces an isomorphism $\mathfrak{h}^{i}(A)\rightarrow\mathfrak{h}^{2g-i}(A)(g-i)$ for all $0\leq i\leq g$. This induces a Lefschetz decomposition of the Chow–Künneth components $\mathfrak{h}^{i}(A)$, see [30, Theorem 5.1]. We are chiefly interested in the components $\mathfrak{h}^{3}(A)$ and $\mathfrak{h}^{2g-3}(A)$ when $g\geq 2$; in that case the Lefschetz decomposition has the form $\mathfrak{h}^{3}(A)=\mathfrak{h}^{3}_{\mathrm{prim}}(A)\oplus\ell\cdot\mathfrak{h}^{1}(A)$ and $\mathfrak{h}^{2g-3}(A)=\mathfrak{h}^{2g-3}_{\mathrm{prim}}(A)\oplus\ell^{g-2}\cdot\mathfrak{h}^{1}(A)$. It has the property that the isomorphism $\ell^{g-3}\colon\mathfrak{h}^{3}(A)\rightarrow\mathfrak{h}^{2g-3}(A)(g-3)$ is a direct sum of isomorphisms $\mathfrak{h}^{3}_{\mathrm{prim}}(A)\rightarrow\mathfrak{h}^{2g-3}_{\mathrm{prim}}(A)$ and $\ell\cdot\mathfrak{h}^{1}(A)\rightarrow\ell^{g-2}\mathfrak{h}^{1}(A)$. Taking Chow groups, we get a decomposition ${\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(A))={\rm CH}^{g-1}(\mathfrak{h}_{\mathrm{prim}}^{2g-3}(A))\oplus{\rm CH}^{g-1}(\ell^{g-2}\cdot\mathfrak{h}^{1}(A))$.
@@ 2.6. Beauville components of $C$
Here we view $[0]-[x_{e}]$ as an element of ${\rm CH}_{0}(J)$.
@@
By definition and the discussion in §2.4, $[C]_{1}\in{\rm CH}_{(1)}^{g-1}(J)={\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J))$. Since $x_{e}$ is torsion, $(n)_{*}([0]-[x_{e}])=0$ for some integer $n\geq 1$. The decomposition (2.2) implies that $[0]=[x_{e}]$ in ${\rm CH}_{0}(J)_{{\rm\mathbb{Q}}}$. Therefore (2.3) shows that $[C]_{0}^{\star(g-2)}\star[C]_{1}=0$. Using properties of the $\mathfrak{sl}_{2}$-action on ${\rm CH}^{*}(J)_{{\rm\mathbb{Q}}}$ (in the sense of [37, §1.3]), this implies that $\ell\cdot[C]_{1}=0$. Since ${\rm CH}^{g-1}(\mathfrak{h}^{2g-3}_{\mathrm{prim}}(J))$ equals the kernel of $\ell\cdot(-)\colon{\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J))\rightarrow{\rm CH}^{g-1}(\mathfrak{h}^{2g-1}(J)(1))$, the corollary follows. ∎
@@ 2.7. Ceresa cycles
Using the expression (2.4), it suffices to prove $[C]_{1}=0$ implies $[C]_{s}=0$ for all $s\geq 1$. This follows from the third centered equation of [51, Theorem 1.5.5]. The proof for $\bar{\kappa}(C)$ is identical. ∎
@@ 3. Vanishing criteria in the Chow and Griffiths groups
By the hypothesis and the Lefschetz isomorphism, $\operatorname{H}^{*}(\mathfrak{h}_{\mathrm{prim}}^{2g-3}(J)^{G})=\operatorname{H}^{2g-3}_{\mathrm{prim}}(J)^{G}=0$. Since any summand of the motive of an abelian variety is finite-dimensional in the sense of Kimura [29, Example 9.1], it follows from Kimura’s [29, Corollary 7.3] that $\mathfrak{h}^{2g-3}_{\mathrm{prim}}(J)^{G}=0$ and hence $[C]_{1}=0$ in ${\rm CH}^{g-1}(J)_{{\rm\mathbb{Q}}}$. By Proposition 2.12, we conclude that $\kappa(C)=0$. ∎
@@
The genus $4$ curve $y^{3}=x^{5}+1$ satisfies $\operatorname{H}^{3}(J)^{{\rm Aut}(C)}=0$ [33, proof of Theorem 3.3], so $\kappa(C)=0$.
@@
We now show that the Hodge conjecture for $J\times A$ implies the claims of the final sentence. Fix mutually inverse isomorphisms of Hodge structures $\phi\colon\operatorname{H}^{3}(J)^{G}\rightarrow\operatorname{H}^{1}(A)(-1)$ and $\psi\colon\operatorname{H}^{1}(A)(-1)\rightarrow\operatorname{H}^{3}(J)^{G}$. By the Hodge conjecture, there exist morphisms of motives $\Phi\colon\mathfrak{h}^{3}(J)^{G}\rightarrow\mathfrak{h}^{1}(A)(-1)$ and $\Psi\colon\mathfrak{h}^{1}(A)(-1)\rightarrow\mathfrak{h}^{3}(J)^{G}$ (in other words, cycles on $J\times A$ with certain properties) with $\operatorname{H}^{*}(\Phi)=\phi$ and $\operatorname{H}^{*}(\Psi)=\psi$. Since $\mathfrak{h}^{1}(A)(-1)$ and $\mathfrak{h}^{3}(J)^{G}$ are Kimura finite-dimensional [29, Example 9.1] and $\operatorname{H}^{*}(\Phi)\circ\operatorname{H}^{*}(\Psi)$ and $\operatorname{H}^{*}(\Psi)\circ\operatorname{H}^{*}(\Phi)$ are the identity, it follows from [29, Proposition 7.2(ii)] (see also [4, Corollaire 3.16]) that $\Psi\circ\Phi$ and $\Phi\circ\Psi$ are themselves isomorphisms. Hence $\Phi$ and $\Psi$ are isomorphisms too and we conclude that $\mathfrak{h}^{3}(J)^{G}\simeq\mathfrak{h}^{1}(A)(-1)$.
Using the Lefschetz isomorphism $\mathfrak{h}^{2g-3}(J)\simeq\mathfrak{h}^{3}(J)(-g+3)$ of [30, Theorem 5.2], we obtain an isomorphism $\mathfrak{h}^{2g-3}(J)^{G}\simeq\mathfrak{h}^{1}(A)(-g+2)$. Similarly to the proof of Theorem A, we decompose $[\iota_{e}(C)]=\sum_{s=0}^{g-1}[C]_{s}$ with $[C]_{s}\in{\rm CH}^{g-1}_{(s)}(J)$ and observe that the class $[C]_{1}$ lies in ${\rm CH}_{(1)}^{g-1}(J)^{G}={\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J)^{G})$. Lemma 2.6 combined with the isomorphism $\mathfrak{h}^{2g-3}(J)^{G}\simeq\mathfrak{h}^{1}(A)(-g+2)$ shows that ${\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J)^{G})={\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J)^{G})_{\mathrm{alg}}$, hence every element of ${\rm CH}^{g-1}_{(1)}(J)^{G}$ lies in ${\rm CH}^{g-1}(J)_{\mathrm{alg},{\rm\mathbb{Q}}}$. Therefore $[C]_{1}\in{\rm CH}^{g-1}(J)_{\mathrm{alg},{\rm\mathbb{Q}}}$, hence the image of $[C]_{1}$ in ${\rm Gr}_{1}(J)_{{\rm\mathbb{Q}}}$ vanishes. Proposition 2.12 then implies that $\bar{\kappa}(C)$ vanishes too. ∎
@@
We now describe some families of curves for which the criterion of Theorem B applies. If $C/k$ is a curve we write $V(C)=\operatorname{H}^{0}(C,\Omega_{C})$, which we think of as an ${\rm Aut}(C)$-representation. Recall that it is $\bigwedge^{3}V(C)\simeq\operatorname{H}^{0}(J,\Omega_{J}^{3})$ that appears in Theorem B.
@@
Let $C\colon y^{3}=x^{4}+ax^{2}+bx+c$ be a Picard curve. Then $V(C)\simeq\chi\oplus\chi\oplus\chi^{2}$ as $C_{3}$-representations, where $\chi$ is a character of order $3$. It follows that $(\bigwedge^{3}V(C))^{{\rm Aut}(C)}=0$, so that the condition of Theorem B is satisfied. We consider these curves in detail in the next section.